# 22affiliationtext: Booth School of Business

### v1

@@ C.1 Evaluation metrics
  • Validation ELBO loss: This is the standard ELBO loss used for both training and validating VAEs. We report this in bits per dimension (bpd), which is the loss scaled by a factor of $\frac{1}{\mathrm{num\_pixels}\cdot\ln 2}$.
  • Reconstruction loss: This approximates the negative log-likelihood. We use binary cross-entropy and report bpd.
  • Covariance error: Unlike the losses above, which evaluate individual reconstructed images with respect to their corresponding ground truth images, this metric (and the ones below) evaluates a reconstructed distribution with respect to a ground truth distribution. Specifically, denoting the empirical covariance matrix across the ground truth datasets as $S^{T}$ and that over the reconstructed dataset as $S^{R}$, we report $\frac{1}{d^{2}}\sum_{i,j=1}^{d}|S^{T}_{ij}-S^{R}_{ij}|$.
  • Mean error: This is analogous to the previous metric but using sample means instead of covariances:$\frac{1}{d}\sum_{i=1}^{d}|\mu^{T}_{i}-\mu^{R}_{i}|$.
  • Maximum mean discrepancy (mmd): Another metric comparing distributions, MMD (Gretton et al., 2012) is a kernel-based metric for comparing two general distributions that may not be distinguishable by merely mean and covariance.
When reporting values for these metrics in tables, we show the mean, averaged over different runs/seeds, followed in parentheses by the standard error. We generally use a 70/30 training/validation split. When appropriate, evaluations are split up so that metrics are computed with respect to the observational context (obs), or averaged over single-concept interventional contexts (ivn), or averaged over double-concept interventional contexts (ood)—with the exception of Table 3, which is detailed in Section 5.2 and Appendix C.4.
C.2 Datasets
In addition to the quad benchmark introduced in Appendix B, we used the following benchmark datasets.
@@ C.2 Datasets
3DIdent
CelebA
We use the original CelebA data (Liu et al., 2015), using the attributes included in the original dataset to define different contexts. Since we do not have access to interventions on this dataset, we relied on conditioning to create different contexts for different concepts. To minimize leakage between contexts, we chose 6 attributes by minimizing the between-attribute correlations and also choosing attributes that are easily detected visually (e.g. black or blond hair). This resulted in six total contexts of varying sizes, one observational along with single-node interventions on bags_under_eyes (41 446 samples), black_hair (48 472 samples), blond_hair (29 983 samples), male (84 434 samples), and mouth_slightly_open (97 942 samples).
C.3 Architectures
We tested our context module using two black-box models:
@@ C.3 Architectures
    lightweight-VAE: A standard 3-layer convolutional network with latent dimension $\dim(\mathbf{z})=128$. This is used for the MNIST and quad experiments in Section 5.1.
@@
    NVAE: A deep hierarchical VAE; see Vahdat and Kautz (2020) for details. This is used for the 3DIdent and CelebA experiments, as well as some MNIST experiments, in Section 5.2.
The lightweight-VAE was used for comprehensive ablations and experiments on simple datasets (150 total models overall, 5 per column per table in Appendix D.1). Since NVAE requires substantially more resources and time to train, this was used to train 6 total models, 2 per dataset.
To provide an apples-to-apples baseline, we trained MNIST using both lightweight-VAE and NVAE.
C.4 Discussion of experimental results (Section 5 in main paper)
Table 1 and Table 2 are interpreted directly as described in Appendix C.1. However, Table 3 warrants some clarification.
Figures D.2.1-D.2.4 in Appendix D.2, show the results of NVAE+context-module across three different datasets (MNIST, 3DIdent, and CelebA). All images here are generated by the model: that is, we draw a sample from the prior and pass it through the context module and decoder—this is to be contrasted with reconstruction, in which a reference image is passed through the encoder to obtain a sample from the posterior that is then passed through the (context module and) decoder. The top row is the result of asking the model to generate images from the observational context (in which no concepts are intervened or conditioned on), while the second and third rows contain images generated from respective contexts in which one concept has been intervened on (in the case of MNIST and 3DIdent) or conditioned on (in CelebA). These first three rows correspond to contexts the model has seen during training (though this is a generation and not reconstruction task). In contrast, the bottom row contains images resulting from asking the model to compose the concepts it has learned (specifically, the two concepts that were individually intervened/conditioned on in the second and third rows)—importantly, notice that the model has not seen data from this double-concept intervention (in the case of MNIST and 3DIdent).
Table 3 compares two different models (namely, NVAE versus NVAE+context-module) across the three datasets. The different models are indicated by the Type column, with “Observational” denoting the plain NVAE model trained on a dataset in which all contexts are pooled, while “Interventional” and “Conditional” denote NVAE+context-module trained on unpooled datasets containing different contexts. The BPD column reports the reconstruction loss (described in Appendix C.1) in standard units so that it is directly comparable with the results of Vahdat and Kautz (2020) and other papers.
C.5 Additional ablations
@@
On MNIST and quad, we ran ablation studies on lightweight-VAE+context-module (denoted “Context module” in tables)—this is our module attached to the black-box, making full use of the different contexts. To evaluate the proposed module, we compared its performance to three natural baselines:
@@
    Ablation 1: lightweight-VAE (without the context module) trained only on the observational context—this is the black-box when denied all (non-observational) context information and data.
@@
    Ablation 2: lightweight-VAE (without the context module) trained on the full pooled dataset—this is the black-box when given all data but no explicit context information.
@@
    Ablation 3 is lightweight-VAE+context-module trained on the full pooled dataset—this is our module attached to the black-box given all the data but not making use of context-specific information.
Note that due to differences in the size of the pooled vs. observational-only datasets, loss values from Ablation 1 are not directly comparable to Ablations 2 or 3 (or the context module), but are included anyway for reference. Results are shown in Section 5.1 and Appendix D.1.1.
@@
We run lightweight-VAE+context-module on MNIST, independently trying out two different regularizers, varying the regularization weight $\lambda$ for each: group lasso and $\ell_{2}$. These are applied to the weights of the reduced-form SEM $A_{0}$ in Equation 1. We also try varying $\beta$ (as in $\beta$-VAE). Results are shown in Appendix D.1.3.
@@
We run lightweight-VAE+context-module on MNIST, jointly varying three parameters that control expressiveness of our context module: expressive input width $0pt_{\mathrm{exp}}$ (which controls latent dimension $\dim(\mathbf{z})$), depth of the expressive layer $h_{\mathrm{exp}}$, and concept width ($0pt_{c}=\dim(\mathbf{c}_{j})=0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ as in Section 4.2). Results are shown in Appendix D.1.4.
Appendix D Additional results
@@
In this section, we present all additional results, including experimental settings described in Appendix C for our context module with the lightweight-VAE black-box architecture as well as additional results on MNIST, 3DIdent, and CelebA for our context module with NVAE.
D.1 lightweight-VAE
D.1.1 Ablations (MNIST and quad)
A description of the architecture and experiment is given in Appendix C. Table 4 shows ablation results after 200 epochs of training on MNIST, and Table 5 is after 1000 epochs. [^1] All other experiments in Appendix D.1 are accordingly run with no ablation and on 200 epochs. These are comparable, with Table 5 indicating some overfitting. Table 6 shows ablation results on quad after 200 epochs.
Altogether, these ablations studies suggest that the cost of the structured representation learned by the context module is only a slight degradation in some of the numerical metrics, while Figure 7 shows sensible OOD concept composition using our module.
Table 4: Ablation on MNIST (200 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.2945 (0.0091) 0.1832 (0.0007) 0.2362 (0.0053) 0.2886 (0.0103)
Reconstruction Loss (bpd) ↓ 0.2649 (0.0093) 0.1469 (0.0008) 0.1998 (0.0050) 0.2575 (0.0111)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.4243 (0.1312) 0.6468 (0.0135) 1.1859 (0.0576) 1.3886 (0.1374)
Mean Error ↓ 0.0218 (0.0033) 0.0052 (0.0005) 0.0204 (0.0018) 0.0237 (0.0053)
MMD ↓ 0.0092 (0.0027) 0.0000 (0.0000) 0.0028 (0.0022) 0.0124 (0.0032)
Covariance Error (ivn) ↓ 2.3717 (0.2764) 2.1983 (0.2567) 2.0005 (0.2380) 2.4591 (0.2833)
Mean Error (ivn) ↓ 0.0332 (0.0030) 0.0221 (0.0019) 0.0258 (0.0030) 0.0339 (0.0039)
MMD (ivn) ↓ 0.0939 (0.0331) 0.0303 (0.0107) 0.0430 (0.0145) 0.0536 (0.0154)
Generated Distribution Evaluation
MMD ↓ 0.0099 (0.0011) 0.0017 (0.0007) 0.0097 (0.0030) 0.0098 (0.0019)
MMD (ivn) ↓ 0.0298 (0.0069) 0.0142 (0.0038) 0.0183 (0.0034) 0.0188 (0.0034)
Table 5: Ablation on MNIST (1000 epochs)
Metric / Evaluation Context module Ablation 1 Ablation 2 Ablation 3
Validation Metrics
Validation ELBO (bpd) ↓ 0.3150 (0.0085) 0.1889 (0.0020) 0.2393 (0.0087) 0.3145 (0.0116)
Reconstruction Loss (bpd) ↓ 0.2836 (0.0087) 0.1513 (0.0025) 0.2010 (0.0090) 0.2837 (0.0112)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.3057 (0.0635) 0.5542 (0.0061) 1.1227 (0.0940) 1.5646 (0.2203)
Mean Error ↓ 0.0154 (0.0023) 0.0040 (0.0001) 0.0194 (0.0038) 0.0274 (0.0096)
MMD ↓ 0.0089 (0.0010) 0.0000 (0.0000) 0.0022 (0.0014) 0.0172 (0.0040)
Covariance Error (ivn) ↓ 2.2713 (0.2320) 2.2229 (0.2536) 1.9782 (0.2247) 2.6567 (0.2901)
Mean Error (ivn) ↓ 0.0314 (0.0023) 0.0221 (0.0018) 0.0263 (0.0030) 0.0392 (0.0046)
MMD (ivn) ↓ 0.0932 (0.0340) 0.0298 (0.0108) 0.0388 (0.0132) 0.0610 (0.0177)
Generated Distribution Evaluation
MMD ↓ 0.0045 (0.0004) 0.0028 (0.0004) 0.0055 (0.0011) 0.0179 (0.0129)
MMD (ivn) ↓ 0.0211 (0.0058) 0.0156 (0.0037) 0.0159 (0.0036) 0.0231 (0.0034)
Table 6: Ablation on quad
Metric / Evaluation Context module Ablation 2
Validation Metrics
Validation ELBO (bpd) ↓ 5.0031 (0.7180) 3.9357 (0.8504)
Reconstruction Loss (bpd) ↓ 4.9586 (0.7161) 3.8795 (0.8506)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 14.5117 (0.5123) 6.0099 (1.3163)
MAE_Mean ↓ 0.0176 (0.0009) 0.0058 (0.0015)
MMD ↓ 0.0000 (0.0000) 0.0000 (0.0000)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 25.3612 (6.8679) 15.8097 (2.4898)
MAE_Mean ↓ 0.0295 (0.0055) 0.0191 (0.0036)
MMD ↓ 0.0060 (0.0030) 0.0022 (0.0015)
Out-of-Distribution (OOD) Evaluation (Averaged)
MAE_Cov ↓ 43.5994 (2.8442) 36.3404 (3.0105)
MAE_Mean ↓ 0.0570 (0.0036) 0.0440 (0.0037)
MMD ↓ 0.0220 (0.0031) 0.0103 (0.0018)
Generated Distribution Evaluation
MMD ↓ 0.0099 (0.0058) 0.0187 (0.0063)
MMD (ivn) ↓ 0.0149 (0.0032) 0.0275 (0.0038)
MMD (ood) ↓ 0.0273 (0.0021) 0.0297 (0.0019)
D.1.2 Examples of concept composition (quad)

### v2

@@ D.1 Evaluation metrics
  • Validation ELBO loss: This is the standard ELBO loss used for validating VAEs. We report this in bits per dimension (bpd), which is the loss scaled by a factor of $\frac{1}{\mathrm{num\_pixels}\cdot\ln 2}$.
  • $p$-sliced Wasserstein (SW) distance: Unlike the loss above, which evaluates individual reconstructed images with respect to their corresponding ground truth images, this metric (Bonneel et al., 2015) evaluates a generated distribution $\mu_{\textrm{gen}}$ with respect to a ground truth distribution $\mu$. Specifically, we use the implementation of Flamary et al. (2021) to compute a Monte-Carlo approximation of $\mathcal{SWD}_{p}(\mu_{\textrm{gen}},\mu)$, where
    $\mathcal{SWD}_{p}(\mu,\nu)=\underset{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}{\mathbb{E}}\left(\mathcal{W}_{p}^{p}(\theta_{\#}\mu,\theta_{\#}\nu)\right)^{\frac{1}{p}},$ (18)
    and $\theta_{\#}\mu$ stands for the pushforwards of the projection $X\in\mathbb{R}^{d}\mapsto\langle\theta,X\rangle$ and $p=2$.
When reporting values for these metrics in tables, we show the mean, averaged over different runs/seeds, followed by the standard error. We generally use a 70/30 training/validation split.
D.2 Datasets
In addition to the quad benchmark introduced in Appendix C, we used the following benchmark datasets.
3DIdent
@@ D.2 Datasets
D.3 Architectures
We tested our context module end-to-end using two black-box models:
@@ D.3 Architectures
    (lightweight) $\beta$-VAE: A standard 3-layer convolutional network with latent dimension $\dim(\mathbf{z})=128$. This is used for the quad experiments in Section 4 and the quad and MNIST experiments in Section E.1.
@@
    NVAE: A deep hierarchical VAE; see Vahdat and Kautz (2020) for details. This is used for the MNIST and 3DIdent experiments in Sections 4 and E.2.
The lightweight-VAE was used for comprehensive ablations and experiments on simple datasets, totaling more than 350 models overall, with 5 per column per table in Appendix E.1. Since NVAE requires substantially more resources and time to train, this was used to train 12 total models, shown in Table 1, the three CM variations plus three ablations on both MNIST and 3DIdent.
D.4 Additional ablations
@@
We ran ablations over different base VAEs augmented with the context module, denoted CM (end-to-end)—this is our module attached to the black-box, making full use of the different contexts, trained end-to-end. To evaluate the proposed module, we compared its performance to three natural baselines:
@@
    CM pooled: CM (end-to-end) but trained with a single context on the full pooled dataset—this is our module attached to the black-box given all the data but not making use of context-specific information, providing an ablation that circumvents our intervention layer (Section 3.2) while maintaining the same architectural complexity.
@@
    base: Base VAE (without the context module) trained only on a single observational context—this is the black-box when denied all (non-observational) context information and data.
@@
    base pooled: Base VAE (without the context module) trained with a single context on the full pooled dataset—this is the black-box when given all data but no explicit context information.
[…]
@@
We run CM (end-to-end) with $\beta$-VAE base on MNIST, independently trying out two different regularizers, varying the regularization weight $\lambda$ for each: group lasso and $\ell_{2}$. These are applied to the weights of the reduced-form SEM $A_{0}$ in 2. We also try varying $\beta$ (as in $\beta$-VAE). Results are shown in Appendix E.1.3.
@@
We run CM (end-to-end) on MNIST, jointly varying three parameters that control expressiveness of our context module: expressive input width $0pt_{\mathrm{exp}}$ (which controls latent dimension $\dim(\mathbf{z})$), depth of the expressive layer $h_{\mathrm{exp}}$, and concept width ($0pt_{c}=\dim(\mathbf{c}_{j})=0pt_{\varepsilon}=\dim(\boldsymbol{\varepsilon}_{j})$ as in Section 3.2). Results are shown in Appendix E.1.5.
D.5 Additional baselines
In addition to the above base models and ablations, we include comparisons against three disentanglement methods, including the unsupervised $\beta$-TCVAE (Chen et al., 2018), the weakly supervised Ada-GVAE (Locatello et al., 2020), and the supervised TVAE (Ishfaq et al., 2018). We gratefully acknowledge the open source implementation of these within the disent Python package (Michlo, 2021). In addition to comparisons against the quad ablations, these baselines are run on MNIST and 3DIdent in Appendix E.1.1.
Appendix E Additional results
@@
In this section, we present complete results, including experimental settings described in Appendix D for our context module with the lightweight-VAE black-box architecture as well as additional results on MNIST and 3DIdent for our context module with NVAE.
E.1 lightweight-VAE
E.1.1 Ablations (MNIST and quad)
A description of the architecture and experiment is given in Appendix D. Tables 25 are already summarized in the main text in Table 1, but here we use the extra space here to additionally present standard errors and show a breakdown of the results over the different contexts.
Table 2: Evaluation of the context module, ablations, and baselines in terms of reconstruction and concept learning on MNIST.
CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
reconstruction (ELBO BPD)
in-distribution 0.259 ± 0.005 0.235 ± 0.004 0.231 ± 0.003 0.181 ± 0.000 0.214 ± 0.003 0.228 ± 0.002 1.035 ± 0.002 1.036 ± 0.012 1.037 ± 0.001
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.099 ± 0.006 0.085 ± 0.002 0.082 ± 0.001 0.067 ± 0.002 0.085 ± 0.003 0.082 ± 0.003 0.521 ± 0.009 0.502 ± 0.014 0.463 ± 0.020
scaled 0.062 ± 0.002 0.100 ± 0.007 0.109 ± 0.011 0.179 ± 0.006 0.142 ± 0.010 0.165 ± 0.008 0.526 ± 0.017 0.492 ± 0.012 0.490 ± 0.016
shear 0.119 ± 0.006 0.110 ± 0.002 0.110 ± 0.003 0.105 ± 0.002 0.109 ± 0.002 0.109 ± 0.003 0.510 ± 0.013 0.476 ± 0.012 0.488 ± 0.024
shift 0.113 ± 0.005 0.112 ± 0.003 0.105 ± 0.003 0.110 ± 0.004 0.115 ± 0.002 0.116 ± 0.004 0.494 ± 0.014 0.447 ± 0.012 0.501 ± 0.009
swel 0.126 ± 0.007 0.113 ± 0.005 0.115 ± 0.002 0.102 ± 0.003 0.129 ± 0.009 0.114 ± 0.003 0.473 ± 0.011 0.492 ± 0.007 0.481 ± 0.011
thic 0.178 ± 0.007 0.184 ± 0.002 0.193 ± 0.010 0.196 ± 0.008 0.231 ± 0.014 0.207 ± 0.014 0.474 ± 0.008 0.475 ± 0.012 0.472 ± 0.013
thin 0.064 ± 0.004 0.080 ± 0.006 0.092 ± 0.010 0.133 ± 0.011 0.107 ± 0.006 0.126 ± 0.007 0.501 ± 0.014 0.506 ± 0.013 0.504 ± 0.011
concept learning mean 0.109 ± 0.006 0.112 ± 0.004 0.115 ± 0.006 0.127 ± 0.005 0.131 ± 0.007 0.131 ± 0.006 0.500 ± 0.012 0.484 ± 0.012 0.486 ± 0.015
Table 3: Evaluation of the context module, ablations, and baselines in terms of reconstruction, concept learning, and composition on quad (independent).
CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
reconstruction (ELBO BPD)
in-distribution 0.519 ± 0.022 0.518 ± 0.006 0.559 ± 0.011 0.444 ± 0.000 0.446 ± 0.002 0.840 ± 0.072 2.017 ± 0.006 1.973 ± 0.001 2.048 ± 0.008
out-of-distribution 0.720 ± 0.069 0.675 ± 0.039 0.774 ± 0.039 3.095 ± 0.295 0.487 ± 0.009 0.830 ± 0.042 2.092 ± 0.007 1.980 ± 0.003 2.115 ± 0.019
concept learning (sliced Wasserstein)
obs 0.056 ± 0.003 0.051 ± 0.004 0.078 ± 0.005 0.061 ± 0.004 0.100 ± 0.007 0.110 ± 0.013 0.332 ± 0.007 0.317 ± 0.009 0.349 ± 0.007
quad1 0.062 ± 0.003 0.056 ± 0.005 0.075 ± 0.003 0.233 ± 0.004 0.201 ± 0.014 0.266 ± 0.020 0.342 ± 0.009 0.335 ± 0.012 0.324 ± 0.013
quad2 0.061 ± 0.005 0.058 ± 0.007 0.086 ± 0.009 0.234 ± 0.014 0.196 ± 0.012 0.254 ± 0.017 0.346 ± 0.007 0.323 ± 0.006 0.334 ± 0.010
quad3 0.063 ± 0.002 0.064 ± 0.005 0.085 ± 0.006 0.251 ± 0.013 0.210 ± 0.008 0.244 ± 0.020 0.333 ± 0.013 0.343 ± 0.010 0.334 ± 0.011
quad4 0.065 ± 0.007 0.057 ± 0.004 0.100 ± 0.011 0.227 ± 0.010 0.227 ± 0.005 0.246 ± 0.011 0.347 ± 0.016 0.334 ± 0.011 0.324 ± 0.016
size 0.075 ± 0.007 0.073 ± 0.006 0.101 ± 0.008 0.157 ± 0.009 0.122 ± 0.001 0.177 ± 0.020 0.327 ± 0.004 0.307 ± 0.010 0.336 ± 0.009
orientation 0.057 ± 0.003 0.053 ± 0.003 0.077 ± 0.006 0.060 ± 0.003 0.104 ± 0.009 0.114 ± 0.016 0.316 ± 0.011 0.349 ± 0.013 0.334 ± 0.010
concept learning mean 0.063 ± 0.004 0.059 ± 0.005 0.086 ± 0.007 0.175 ± 0.008 0.166 ± 0.008 0.202 ± 0.017 0.335 ± 0.010 0.330 ± 0.010 0.334 ± 0.011
composition (sliced Wasserstein)
(quad1, quad2) 0.100 ± 0.028 0.108 ± 0.016 0.137 ± 0.014 0.311 ± 0.009 0.276 ± 0.014 0.343 ± 0.018 0.346 ± 0.012 0.346 ± 0.004 0.334 ± 0.007
(quad1, quad3) 0.078 ± 0.016 0.107 ± 0.011 0.146 ± 0.013 0.323 ± 0.011 0.273 ± 0.007 0.349 ± 0.008 0.367 ± 0.016 0.372 ± 0.009 0.337 ± 0.008
(quad1, quad4) 0.094 ± 0.032 0.091 ± 0.010 0.179 ± 0.014 0.346 ± 0.017 0.277 ± 0.017 0.325 ± 0.023 0.377 ± 0.018 0.343 ± 0.012 0.346 ± 0.010
(quad1, size) 0.101 ± 0.010 0.087 ± 0.005 0.111 ± 0.010 0.266 ± 0.012 0.226 ± 0.013 0.284 ± 0.011 0.343 ± 0.005 0.331 ± 0.009 0.336 ± 0.009
(quad1, orientation) 0.060 ± 0.003 0.059 ± 0.004 0.080 ± 0.002 0.236 ± 0.012 0.194 ± 0.017 0.240 ± 0.004 0.367 ± 0.016 0.360 ± 0.007 0.356 ± 0.007
(quad2, quad3) 0.067 ± 0.007 0.098 ± 0.017 0.148 ± 0.014 0.338 ± 0.017 0.283 ± 0.014 0.325 ± 0.016 0.350 ± 0.010 0.349 ± 0.012 0.350 ± 0.008
(quad2, quad4) 0.068 ± 0.004 0.078 ± 0.010 0.194 ± 0.018 0.302 ± 0.013 0.291 ± 0.015 0.332 ± 0.012 0.382 ± 0.007 0.342 ± 0.005 0.340 ± 0.013
(quad2, size) 0.099 ± 0.008 0.083 ± 0.008 0.102 ± 0.007 0.248 ± 0.008 0.197 ± 0.004 0.281 ± 0.021 0.334 ± 0.007 0.328 ± 0.015 0.336 ± 0.006
(quad2, orientation) 0.063 ± 0.004 0.058 ± 0.003 0.090 ± 0.012 0.225 ± 0.011 0.214 ± 0.010 0.274 ± 0.013 0.329 ± 0.011 0.348 ± 0.010 0.333 ± 0.008
composition mean 0.081 ± 0.013 0.086 ± 0.009 0.132 ± 0.012 0.288 ± 0.012 0.248 ± 0.012 0.306 ± 0.014 0.355 ± 0.011 0.346 ± 0.009 0.341 ± 0.008
Table 4: Evaluation of the context module, ablations, and baselines in terms of reconstruction, concept learning, and composition on quad (dependent).
CM (end-to-end) CM (fine-tune) CM (frozen) base base pooled CM pooled Ada-GVAE BetaTCVAE TVAE
reconstruction (ELBO BPD)
in-distribution 0.828 ± 0.004 0.840 ± 0.005 0.889 ± 0.017 0.767 ± 0.001 0.746 ± 0.000 0.752 ± 0.000 2.443 ± 0.002 2.408 ± 0.000 2.455 ± 0.002
out-of-distribution 0.863 ± 0.006 0.908 ± 0.005 0.959 ± 0.011 0.815 ± 0.005 0.746 ± 0.001 0.753 ± 0.001 2.450 ± 0.002 2.406 ± 0.000 2.463 ± 0.002
concept learning (sliced Wasserstein)
obs 0.107 ± 0.005 0.103 ± 0.008 0.116 ± 0.003 0.110 ± 0.005 0.104 ± 0.006 0.113 ± 0.004 0.365 ± 0.007 0.375 ± 0.008 0.357 ± 0.005
quad1 0.100 ± 0.004 0.102 ± 0.004 0.107 ± 0.008 0.158 ± 0.003 0.163 ± 0.004 0.137 ± 0.007 0.373 ± 0.004 0.352 ± 0.004 0.374 ± 0.008
quad2 0.103 ± 0.004 0.100 ± 0.006 0.110 ± 0.009 0.179 ± 0.007 0.165 ± 0.008 0.173 ± 0.009 0.373 ± 0.004 0.359 ± 0.003 0.366 ± 0.005
quad3 0.127 ± 0.017 0.127 ± 0.008 0.139 ± 0.008 0.200 ± 0.010 0.214 ± 0.009 0.213 ± 0.015 0.379 ± 0.005 0.374 ± 0.006 0.372 ± 0.009
quad4 0.145 ± 0.010 0.140 ± 0.007 0.131 ± 0.011 0.305 ± 0.012 0.277 ± 0.010 0.278 ± 0.015 0.401 ± 0.015 0.386 ± 0.008 0.388 ± 0.013
size 0.109 ± 0.006 0.105 ± 0.006 0.117 ± 0.004 0.116 ± 0.004 0.121 ± 0.007 0.119 ± 0.004 0.361 ± 0.006 0.362 ± 0.005 0.367 ± 0.005
orientation 0.123 ± 0.003 0.109 ± 0.007 0.124 ± 0.004 0.120 ± 0.009 0.123 ± 0.008 0.130 ± 0.007 0.369 ± 0.006 0.353 ± 0.002 0.365 ± 0.007
concept learning mean 0.116 ± 0.007 0.112 ± 0.007 0.121 ± 0.007 0.170 ± 0.007 0.167 ± 0.007 0.166 ± 0.009 0.375 ± 0.007 0.366 ± 0.005 0.370 ± 0.007
composition (sliced Wasserstein)
(quad1, quad2) 0.111 ± 0.008 0.097 ± 0.007 0.123 ± 0.011 0.229 ± 0.003 0.214 ± 0.005 0.196 ± 0.013 0.366 ± 0.006 0.381 ± 0.010 0.367 ± 0.007
(quad1, quad3) 0.134 ± 0.008 0.117 ± 0.003 0.140 ± 0.008 0.227 ± 0.009 0.228 ± 0.008 0.198 ± 0.011 0.390 ± 0.009 0.367 ± 0.011 0.374 ± 0.004
(quad1, quad4) 0.133 ± 0.007 0.163 ± 0.010 0.163 ± 0.008 0.284 ± 0.015 0.292 ± 0.006 0.276 ± 0.013 0.403 ± 0.012 0.376 ± 0.013 0.361 ± 0.006
(quad1, size) 0.108 ± 0.003 0.106 ± 0.004 0.123 ± 0.007 0.172 ± 0.009 0.170 ± 0.009 0.157 ± 0.011 0.360 ± 0.008 0.357 ± 0.008 0.367 ± 0.003
(quad1, orientation) 0.110 ± 0.007 0.108 ± 0.007 0.125 ± 0.006 0.167 ± 0.006 0.157 ± 0.007 0.143 ± 0.007 0.365 ± 0.010 0.371 ± 0.007 0.369 ± 0.004
(quad2, quad3) 0.158 ± 0.007 0.150 ± 0.013 0.164 ± 0.005 0.222 ± 0.008 0.200 ± 0.011 0.195 ± 0.010 0.367 ± 0.007 0.378 ± 0.006 0.379 ± 0.008
(quad2, quad4) 0.130 ± 0.011 0.122 ± 0.009 0.119 ± 0.010 0.299 ± 0.008 0.282 ± 0.025 0.279 ± 0.008 0.403 ± 0.010 0.404 ± 0.013 0.382 ± 0.014
(quad2, size) 0.126 ± 0.009 0.109 ± 0.006 0.120 ± 0.005 0.191 ± 0.011 0.166 ± 0.010 0.181 ± 0.010 0.368 ± 0.004 0.365 ± 0.005 0.363 ± 0.004
(quad2, orientation) 0.135 ± 0.004 0.120 ± 0.010 0.133 ± 0.006 0.208 ± 0.011 0.195 ± 0.020 0.198 ± 0.010 0.368 ± 0.005 0.366 ± 0.005 0.351 ± 0.006
composition mean 0.127 ± 0.007 0.121 ± 0.008 0.134 ± 0.007 0.222 ± 0.009 0.212 ± 0.011 0.203 ± 0.010 0.377 ± 0.008 0.374 ± 0.009 0.368 ± 0.006
Table 5: Evaluation of the baselines in terms of reconstruction, concept learning, and composition on 3DIdent.
Ada-GVAE BetaTCVAE TVAE
reconstruction (ELBO BPD)
in-distribution 3.267 ± 0.001 3.261 ± 0.000 3.282 ± 0.005
out-of-distribution 3.274 ± 0.001 3.264 ± 0.001 3.291 ± 0.006
concept learning (sliced Wasserstein)
obs 0.267 ± 0.011 0.257 ± 0.012 0.245 ± 0.003
bg 0.232 ± 0.012 0.254 ± 0.009 0.262 ± 0.006
obj 0.243 ± 0.007 0.263 ± 0.010 0.259 ± 0.007
sl 0.253 ± 0.013 0.254 ± 0.007 0.250 ± 0.009
concept learning mean 0.249 ± 0.011 0.257 ± 0.010 0.254 ± 0.006
composition (sliced Wasserstein)
(bg, obj) 0.252 ± 0.014 0.261 ± 0.011 0.256 ± 0.010
(bg, sl) 0.227 ± 0.009 0.255 ± 0.012 0.241 ± 0.013
(obj, sl) 0.274 ± 0.010 0.247 ± 0.005 0.269 ± 0.009
composition mean 0.251 ± 0.011 0.254 ± 0.009 0.256 ± 0.011
E.1.2 Examples of concept composition (quad)
Figure LABEL:fig:quad-ood-full depicts samples generated from the context module appended to the lightweight-VAE described in Appendix D.3.
[…]
fig:quad-ood-full